OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (ribhu):

Multiply the first equation by 2

OpenStudy (anonymous):

ok

OpenStudy (ribhu):

then subtract the 2nd from 1st equation

OpenStudy (anonymous):

wait would you plug in 2 for one of the variables

OpenStudy (ribhu):

you will get x, and substituting the value of x in any of those equation you will get y, hence the question is solved

OpenStudy (ribhu):

10x+2y=18 3x+2y=4 subtracting we get 7x=14 x=2

OpenStudy (ribhu):

putting x=2 in 2nd equation we get 3(2)+2y=4 y=-1

OpenStudy (anonymous):

oh thank you but I just solved it by plugging the options given to me but thanks because that will help on non multiple choice test

OpenStudy (ribhu):

ok